-What is the boiling point in oC of a 0.30 M solution of NaHSO4 if the solution has a density of 1.002 g/mL. Ka for hydrogen sulfate ion is 1.2 x 10–2?
Answer #1

Molarity = 0.3M dersity = 1.002 gmime 0.3 mol in I litre. » V=1 litr = 1000 me MM NaHSO4 = 23 +1 +32+(4+16) d: 1.002 gm/me maОть = ixk, Уm - 2-2 o-S12. x 0 - 3 05 6 Дт. = 03yer с Tь - Тъ = озчіс с - 100 = с. 35 - 109-352)